A 31 g sample of Cl2 (Chlorine) gas has a volume of 521 mL at a temperature of 60°C.
i. Calculate the pressure using the Ideal Gas Law.
ii. Calculate the corrected pressure using the Van der Waals Equation.
atm×L2 L a=6.254 mol2 b=0.05422 mol
1
Expert's answer
2020-04-15T01:51:16-0400

T=60+273=333 K

Mr(Cl2) = 35.5*2=71

i. Ideal Gas Law : piVi=m(Cl2)/Mr(Cl2) *RT

pi=m(Cl2)*R*T/(Mr(Cl2) *Vi)=(31*8.314*333) / (71*0.521)=2320.17 kPa

ii. Van der Waals Equation: [ pr + a*(n/V)2](Vr - n*b)=nRT

pr= (nRT)/(V-n*b) - a*(n/V)2 = ((31/71)*8.314*333)/(0.521 - (31/71)*0.05422) - 6.254*((31/71)/0.521)^2)=2426.22 kPa